OpenStudy (anonymous):

What is the product written in scientific notation? (4.5 * 10^-8) ( 6*10^-4) A.2.7 * 10^32 B. 2.7 * 10^-13 C. 2.7 * 10^-12 D. 2.7 * 10^-11

OpenStudy (welshfella):

fist multiply the first parts of the numbers 4.5 * 6 = ?

OpenStudy (welshfella):

what does that come to?

OpenStudy (anonymous):

27

OpenStudy (welshfella):

right

OpenStudy (welshfella):

now multiply 10/6-8 * 10^-4 add the exponents -8 and -4

OpenStudy (welshfella):

* 10^-8 * 10*-4

OpenStudy (anonymous):

-12

OpenStudy (welshfella):

ok that gives us 27 * 10^-12 but this is not scientific notation the first part must have a value between 1 and 10

OpenStudy (welshfella):

so we divide the first part by 10 and , in order to maintain the correct value, multiply the exponent part by 10

OpenStudy (anonymous):

Add a decimal?

OpenStudy (anonymous):

To get 2.7

OpenStudy (welshfella):

2.7 / 10 = ? 10^-12 * 10^1 = ?

OpenStudy (welshfella):

yes 2.7 is correct for first part

OpenStudy (anonymous):

2.7 / 10 = 0.27

OpenStudy (welshfella):

10^-12 * 10^1 we ADD the exponents 10^-12 * 10^1 = 10^(-12+1) = ?

OpenStudy (anonymous):

-11

OpenStudy (welshfella):

No 27 / 10 = 2.7 is correct for first part (2.7 is between 1 and 10)

OpenStudy (welshfella):

-11 is right so what is the correct option?

OpenStudy (anonymous):

D

OpenStudy (welshfella):

yep

OpenStudy (anonymous):

Thank you! ^-^/

OpenStudy (welshfella):

yw

OpenStudy (welshfella):

general form for scientific notation is N * 10^n where N is a number between 1 and 10 ( not including 10)